NottheMonga2087:

Solve the system of equations by substitution. 3/8x+1/3y=17/24 x+7y=8 ( ) ( )

supie:

@az 👀

xxDeppressionxx:

[1] 3x/8 + y/3 = 17/24 [2] x + 7y = 8 To remove fractions, multiply equations by their respective LCD Multiply equation [1] by 24 Equations now take the shape: [1] 9x + 8y = 17 [2] x + 7y = 8 Solve equation [2] for the variable x [2] x = -7y + 8 Plug this in for variable x in equation [1] [1] 9•(-7y+8) + 8y = 17 [1] - 55y = -55 Solve equation [1] for the variable y [1] 55y = 55 [1] y = 1 By now we know this much : x = -7y+8 y = 1 Use the y value to solve for x x = -7(1)+8 = 1 solution {x,y} = {1,1}

xxDeppressionxx:

Source is (www.tiger-algebra.com)

NottheMonga2087:

thank youuu

xxDeppressionxx:

Np